The water-gas shift reaction is an important source of hydrogen. The value of Kc for the reaction below is is 5.1 at 700 K. CO+Hg) Co(8)+H2(8 ﹀ 1st attempt Part 1 (1point) See Periodic Table Calculate t the equilbrium concentrations of the four gasesif the initial concentration of each of them is 00954 M Calculate the equilibrium concentrations of the four gases if the initial concentration of each of them is 0.0954 M. col Part 2 (1 point) CO21
Answer #1

The reaction given

CO(g) + H2O(g) --------> CO2(g) + H2(g)

Kc= 5.1

it is given that initially 0.0954 M of each of them is present

CO(g)    +           H2O(g) -------->            CO2(g)     +    H2(g)

0.0954                      0.0954                         0.0954               0.0954

0.0954-x                 0.0954-x                         0.0954+x           0.094+x

Kc expression will be

Keーで에H201

0.0954] 0.0954+ 0.0954- 0.0954-

    5.1 [0.0954-1]*[0.0954-x] = [0.0954+x] * [0.0954+x]

5.1 (0.0954-x)2 = (0.0954+x)2

5.1 ( 0.09542 - 2 * 0.0954*x + x2 ) = ( 0.09542 + 2 * 0.0954*x + x2 )

5.1* 0.09542 -5.1* 2 * 0.0954*x+ 5.1x2 = 0.09542 + 2 * 0.0954*x + x2

0.046416 - 0.97308*x + 5.1 x2   = 9.10116*10-3 +0.1908*x+x2

( 5.1-1)x2 -( 0.97308+0.1908)*x +0.046416-9.10116*10-3 =0

4.1 x2 -1.16388x+0.03731484 = 0

on solving the above quadratic equation , x= 0.24703, 0.0368423

Since the initial concentration was 0.0950 , 0.247 cannot be root as that would been negative molar concentration of reactant at equilibrium which is not possible

Hence , x =0.0368423 M

Hence

PART (1)

[CO(g) ]= 0.0950-x=0.0950-0.0368423 = 0.0581577 M

PART (2)

[CO2(g) ]= 0.0950+x= 0.0950+0.0368423 = 0.1318423 M

PART (3)

[H2O(g)] =0.0950 -x = 0.0950-0.0368423 =0.0581577 M

PART (4)

[H2(g) ]=0.0950+x = 0.0950+0.0368423 = 0.1318423 M